# PathoSyn: Imaging-Pathology MRI Synthesis via Disentangled Deviation Diffusion

### v1

@@ 5 Discussion & Conclusion
The PathoSyn framework advances medical image synthesis by shifting from unstructured holistic generation to a principled decomposition of pathology. By modeling a stochastic deviation field over a deterministic anatomical substrate, it directly addresses the trade-off between anatomical preservation and pathological realism. Quantitative analysis shows that the clinical utility of synthetic data depends on its distributional and structural consistency. The low feature correlation (C=0.072) acts as a safeguard against anatomical leakage, ensuring that the patient’s underlying identity remains intact during disease synthesis. By capturing the high-frequency textural heterogeneity and non-Gaussian intensity fluctuations of high-grade gliomas, PathoSyn reduces shortcut learning and forces downstream models to learn genuine radiomic signatures, such as infiltrative margins and necrotic cores, that transfer to real-world clinical scenarios. Grounded in a diffusion training objective with seam-aware blending and boundary consistency, the framework removes the need for inference-time heuristics and yields seamless, anatomically coherent synthesis, enabling multiple biologically credible digital twins for stress-testing clinical algorithms on rare morphologies. Future work will extend PathoSyn into a fully end-to-end generative pipeline that also models non-lesion structural deformations. In particular, we aim to capture mass effect, where pathological growth induces global geometric shifts and compression of healthy brain tissues, including ventricular displacement and midline shifts. Unifying these photometric and morphometric changes in a single generative system will be key for building the next generation of resilient, accurate, and trustworthy medical AI diagnostic tools.

### v2

@@ 5 Discussion & Conclusion
In this paper, we introduce PathoSyn, a pathology-aware medical image synthesis framework that disentangles stable anatomical structure from stochastic pathological variation. Instead of synthesizing complete images from a single joint distribution, PathoSyn represents disease as a deviation field superimposed on the preserved anatomy. This formulation limits identity leakage, protects the patient’s specific structure, and improves the fidelity of tumor-related appearance changes. Compared to state-of-the-art approaches (holistic diffusion models, mask-based inpainting, and VAE–GANs), PathoSyn produces images with sharper infiltrative tumor margins, more realistic edema and necrotic textures, and fewer structural distortions. The low empirically measured correlation between anatomical and pathological representations (C = 0.072) indicates that pathology can be modulated with minimal unintended alteration of the anatomy of the core brain. Our framework removes the need for manual post hoc blending or heuristic inference-time adjustments, enabling seamless synthetic images for downstream research and clinical assessment. More broadly, the deviation field formulation provides a conceptual tool for modeling structured variation in computer vision and machine learning. It enables the representation of subtle changes in internal features and localized textural shifts without altering the global geometry. This principle extends to other medical imaging modalities such as CT, PET, ultrasound, and digital pathology, where disease progression often appears as regional texture and intensity changes rather than large-scale shape deformation.
A prominent clinical use case for PathoSyn is preoperative planning for brain tumor surgery. By producing multiple plausible MRI variants for an individual patient, PathoSyn can support clinicians in anticipating the short-term evolution of tumor texture, infiltration patterns, and peritumoral edema in the interval preceding surgery. These synthetic scenarios may help neurosurgeons and radiologists in the operative planning, risk stratification, and evaluation of alternative strategies in settings where updated images are unavailable or when rapid disease progression is suspected. However, the current implementation still has potential for further refinement. PathoSyn primarily models appearance changes and does not yet simulate large-scale structural deformations such as mass effect, ventricular compression, or midline shift. In addition, robustness across scanners, institutions, and acquisition protocols remains constrained by domain shift, indicating a need for improved domain adaptation and harmonization. These issues must be resolved before the framework can be considered for regulated clinical deployment.
Future work will extend PathoSyn into a unified generative framework capable of jointly modeling both appearance alterations and geometric deformations. This expansion will enable representation of not only tumor texture but also the spatial impact of neoplastic growth on adjacent brain structures. The overarching objective is to support realistic data generation for surgical planning, longitudinal disease monitoring, and robustness evaluation of machine learning models. By explicitly combining anatomical stability with controlled, realistic pathological variation, PathoSyn constitutes a step toward clinically meaningful synthetic data generation that can underpin more reliable, interpretable, and trustworthy medical AI systems.